A mixture of 2 moles of helium gas (atomic mass = 4 amu) and 1 mole of argon gas (atomic mass = 40 amu) is kept at 300 K in a container. The ratio of the rms speeds

Grade:9

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

v=root(3rt/m)
=>ratio =root(m_argon/m_helium)=root(10)
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
vrms = (3RT/M)
=>ratio = (MAr/MHe) = 10 : 1 = 3.16 : 1
 
Thanks and regards,
Kushagra